Calculate (a) the pH and (b) the concentration of
Answer #1

Henderson-Hasselbalch Equation :

pH = pKa + logfM [HA]

Ka for formic acid is 1.8 * 10^-4 , so pka -logka = 3.75

so from above equation ph = 3.75 + log (0.1/0.1) = 3.75 +log1 , we know log1 =0

ph = 3.75

from above question

N1 = 0.M = 0.1N

V1 = 400ML

N2= 0.1M = O.1 N

V2 = 100 ML

V = 500 ML

N = ?

we know N1V1 + N2V2 = NV

400 * 0.1 + 100 *0.1 = 500 * x

x= 0.1 M

Concentaration of acettae ion is 0.1 M